OpenStudy (anonymous):

plz hlp me to solve this equation y-x²-6x+9=0 XD

OpenStudy (anonymous):

sorry but that doesn't help :(

OpenStudy (anonymous):

y-x²-6x+9=0 y=x²+6x-9=0 or x²+6x-9=0 can you go from here to solve the quadratic eq?

OpenStudy (anonymous):

no i'm supposed to find the expression of the inverse function of this f(x)=(x-3)²-1 i did this f(x)=y <=> f^-1(y)=x <=> y=(x-3)²-1 now i must find the expression by x plz help me

OpenStudy (anonymous):

so this is not the problem anymore?

OpenStudy (anonymous):

y-x²-6x+9=0

OpenStudy (anonymous):

f(x)=(x-3)²-1 then y=(x-3)²-1 y+1=(x-3)² (y+1)^1/2=(x-3) (y+1)^1/2 +3= x now the inverse is f^-1(y)=(y+1)^1/2 +3 ans

OpenStudy (anonymous):

thanx you very much :)

OpenStudy (anonymous):

yw

OpenStudy (anonymous):

good luck now